Use 1,3,5,7,9,11,13,15 and fill the blanks___+___+___=30

Grade:12th pass

3 Answers

CHALLA NAVNEESWAR REDDY
7 years ago
given 1,3,5,7,9,11,13,15     
we want to fill_+_+_=30
from the given numbers
11+2(7)+1
=11+14+1
=30
hence it is proved
SAHIL
7 years ago
ANSWER:
3! + 11 + 13
THIS MEANS 3 FACTORIAL + 11 + 13
WHERE 3 FACTORIAL = 3 X 2=6
AND 6+ 11+ 13 = 30
THANKS
KEEP POSTING
LET ME KNOW IF THERE IS SOME DOUBT
Adarsh
6 years ago
It is not possible to add three odd numbers to get an even number.let that first number be 2n+1,2n+3,2n+5(where n is an even interfere): adding all these gives us 6n+9 =3(2n+3) which is an odd number.. hence the question is wrong